A DNA has 5.6 Kilo base pairs. If no. of cytosine molecule in this gene is 1000 then what will be the no. of thymine molecule in this gene ?

Grade:12

5 Answers

Arun
3 years ago
6.2kilo base pair=6.2 multiply with 1000= 6200 base pair
T+G=6200
2160+G=6200
G= 6200-2160
G= 4040 base pair
cassia
3 years ago
1800 is the number of thymine molecule. but explaination needed.
Ayush_Deep
3 years ago
Total base pair = 5600
Total nucleotide= 11200
Cytosine = guanine = 1000
Cytosine + guanine = 2000
Adenine + thymidine = 9200
Thymidine = adenine = 9200/2= 4600
 
Laksha Koli
3 years ago
total base pairs= 5.6 K = 5600
Total cytosine = 1000 = guanine
C+G= 2000
A+T = 5600 – 2000
        = 3600
2T = 3600
T = 1800
HARSHVARDHAN GAUR
3 years ago
GIVEN THAT 5.6 KILO BASE PAIR ARE GIVEN THEN                                                                                                           5.6*1000 IS EQUAL TO 5600                                                                                                                                             GIVEN THAT THERE ARE 1000 CYTOSINE                                                                                                                          WE KNOW THAT,   NO. OF CYTOSINE IS EQUAL TO NO. OF GUANINE 
SO  ADD CYTOCINE AND GUANINE NO. AND  THAT ARE 1000 AND 1000 THEN WE GET 2000
NOW SUBTRACT 2000 FROM 5600 THEN WE WILL GET 3600 
AND WE KNOW THAT NO. OF ADENINE IS EQUAL TO NO. OF THYMINE 
SO,     DIVIDE 3600 INTO TWO EQUAL PARTS 3600/2  THEN WE GET 1800 
HENCE ADENINE NO. IS EQUAL TO 1800 AND AS SAME AS THYMINE IS ALSO EQUAL TO 1800